An alloy of Au and Ag have Au at corners and Ag at face centres of a cube. Specify the formula of alloy.

Grade:12

1 Answers

Vikas TU
7 years ago
At corners the proportion of each atoms is 1/8.
and face centers the proprotion for each atom is = ½.
therefore Au = (1/8)*8 = 1
and
Ag = (1/2)*6 = 3
formulae:
Au(Ag)3